OpenStudy (anonymous):

Will give medal! Graph Y= 1+1/2 3^sqrt x-2

OpenStudy (phi):

is this \[ y= \frac{1}{2}3^{\sqrt{x-2}} \]?

OpenStudy (anonymous):

\[y= 1 + \frac{ 1 }{ 2 } \sqrt[3]{x-2}\]

OpenStudy (anonymous):

sorry it is like the one i just posted.

OpenStudy (phi):

do you have to do it "by hand" ?

OpenStudy (phi):

or can you use a graphing calculator or geogebra (free download)?

OpenStudy (anonymous):

here are the choices.

OpenStudy (phi):

are you allowed to use a calculator?

OpenStudy (anonymous):

i dont have one.

OpenStudy (anonymous):

but im pretty sure D is wrong

OpenStudy (phi):

one way to do these problems is "test out" some "easy" points do you know how to find the cube root of 8 ? i.e. \( \sqrt[3]{10-2} \)

OpenStudy (anonymous):

no i dont

OpenStudy (phi):

what number times itself 3 times is 8 ? not 1*1*1 =1 how about 2*2*2 ?

OpenStudy (anonymous):

im sorry i found an online graphing calculator. desmos.com i found the answer but i will still give you a medal! Thanks!

OpenStudy (phi):

yes, graphing is one way. But if you have to do this without graphing, it might be worth learning how to do it without a graph.